# Escaping the Subprime Trap in Algorithmic Lending

### v1

@@ Abstract
Disparities in lending to minority applicants persist even as algorithmic lending practices proliferate. Further, disparities in interest rates charged can remain large even when loan applicants from different groups are equally creditworthy. We study the role of risk-management constraints, specifically Value-at-Risk ($\mathrm{VaR}$) constraints, in the persistence of segregation in loan approval decisions. We develop a formal model in which a mainstream (low-interest) bank is more sensitive to variance risk than a subprime (high-interest) bank. If the mainstream bank has an inflated prior belief about the variance of the minority group, it may deny that group credit indefinitely, thus never learning the true risk of lending to that group, while the subprime lender serves this population at higher rates. We formalize this as a “subprime trap” equilibrium. Finally, we show that a small, finite subsidy (or partial guarantee) can help minority groups escape the trap by covering enough of the mainstream bank’s downside so that it can afford to lend and learn the minority group’s true risk. Once it has sufficiently many data points, it meets its VaR requirement with no further assistance, minority groups are approved for loans by the mainstream bank, and competition drives down the interest rates of subprime lenders.
1 Introduction
@@
Algorithmic lending has grown rapidly as scalable ML methods achieve wide adoption among both existing lenders and new market entrants, bringing with them the possibility of significantly improving the fairness of financial decisions based on observable data about loan applicants Khandani et al. [2010], Bono et al. [2021], Berg et al. [2022], Remolina [2022]. However, inequalities persist in both loan approval rates and interest rates charged to minority applicants versus white applicants, and a switch to algorithmic lending procedures does not necessarily improve outcomes on either metric Romei and Ruggieri [2013], Žliobaitė [2017], Quillian et al. [2020], Giacoletti et al. [2021], Aliprantis et al. [2022], Bartlett et al. [2022], Fuster et al. [2021]. These inequities exist against a background of historic discrimination in US retail banking to individuals Quillian et al. [2020] and businesses Blanchflower et al. [2003] and a long-standing “racial wealth gap” – persistent differences in median household wealth by ethnic group Charles and Hurst [2002], Derenoncourt et al. [2023], Althoff and Reichardt [2024].
@@
Notably, disparities in interest rates and loan approvals persist even when minority applicants have comparable credit scores to majority applicants Bayer et al. [2016], Popick [2022], Crosignani and Le [2023]. This combination of facts creates a puzzle: are lenders selecting on observable risk factors? And if not, why, besides explicit discrimination, might they be failing to lend to minority applicants?
@@
Several empirical findings help us to understand this puzzle. First, ethnic groups sort across lenders, with minorities more likely to accept loans from high-interest rate banks versus conventional lenders, even conditional on credit score Bayer et al. [2016]. Bayer et al. [2016] writes:
@@
It is this set of stylized facts that motivates our model. How do minorities with good credit scores end up with no choice but to accept high-cost loans?
@@
To answer this question, we consider the role of risk management constraints when banks have imperfect information about their applicant pool. Lending is a problem of imperfect information: banks do not know the exact probability that a given borrower will repay a loan, and so must use observable characteristics as a proxy for repayment probability Adams et al. [2009], Crawford et al. [2018]. Seminal work from Coate and Loury [1993] shows how inaccurate beliefs can sustain equilibria in which minority groups systematically lose out Coate and Loury [1993], Kim and Loury [2018].
@@
Specifically, we study the effect of Value-at-Risk ($\mathrm{VaR}$) constraints on loan approval decisions Artzner et al. [1999], Jorion [2000]. Our analysis focuses on a formal model in which a mainstream, or “low interest rate”, bank exhibits heightened sensitivity to variance risk relative to a subprime, or “high interest rate”, bank. We show that when the mainstream bank has inflated prior beliefs about the variance repayments from the minority group, $\mathrm{Var}$ constraints bind, which means that it will then systematically refrain from extending credit to that group. This reluctance then prevents the bank from updating its beliefs regarding the true risk, thereby locking minority borrowers into higher-cost subprime lending arrangements. This mechanism is characterized as a “subprime trap”.
1.1 Our Contributions
@@
First, we contribute an analysis of risk management constraints to the study of algorithmic fairness. We help address a puzzle of general interest: why might banks fail to extend loans even when they would otherwise be profitable? Second, we bring details of real-world banking practices into the study of algorithmic fairness. We introduce a two-bank equilibrium framework wherein the mainstream bank enforces a more stringent $\mathrm{VaR}$ requirement than its subprime counterpart. Under these conditions, an erroneous initial estimate of the minority group’s variance leads the mainstream bank to permanently refrain from lending to that group. Second, we formalize the resulting equilibrium (1) and demonstrate that the minority group is consequently confined to subprime loans, despite having identical average creditworthiness as the majority group. Third, we establish that the provision of a modest, finite subsidy or partial guarantee can resolve this equilibrium inefficiency. By mitigating the downside risk sufficiently, the subsidy enables the mainstream bank to lend to and learn about the creditworthiness of the minority group. With an updated, more accurate assessment of risk, the bank eventually meets its $\mathrm{VaR}$ requirements without continued external intervention, thereby extending mainstream credit at favorable rates.
@@
This investigation contributes to the literature by highlighting how inaccurate risk metrics, in conjunction with risk-based capital constraints, can systematically exclude certain borrower groups from low-interest credit. Our results further imply that targeted subsidies may effectively rectify informational failures, yielding improved outcomes for both lenders and minority borrowers.
1.2 Related Literature
@@
Our work contributes to several strands of literature, including algorithmic fairness in lending, the economics of discrimination, and the role of informational frictions in sustaining suboptimal market equilibria. In the domain of algorithmic fairness, recent research has examined how increasingly sophisticated data-driven methods in credit scoring may inadvertently perpetuate or exacerbate disparities Barocas and Selbst [2016], Barocas et al. [2023], Kumar et al. [2022]. While algorithmic approaches promise enhanced accuracy, empirical evidence indicates that minority applicants often face higher interest rates or are denied credit outright Bartlett et al. [2022], Crosignani and Le [2023].
@@
Discrimination is a classical topic in economics research, studied in a large theoretical and empirical literature Becker [1971], Phelps [1972], Lundberg and Startz [1983], Arrow [1998], Heckman [1998], Cowgill and Tucker [2019]. Building on this tradition, subsequent studies have explored how negative stereotypes or misperceptions of risk may translate into adverse outcomes for minority borrowers Fang and Moro [2011], Coate and Loury [1993]. In a manner analogous to these analyses, our work demonstrates that inflated beliefs about repayment variance can result in a self-reinforcing equilibrium whereby minority groups are persistently relegated to subprime lending markets.
A related literature examines the role of informational frictions and the corrective impact of policy interventions Akerlof [1970], Akerlof and Shiller [2015], Thaler [2016]. Authors such as Cai et al. [2020] have investigated contexts in which banks can acquire improved information about borrowers’ creditworthiness through selective experimentation. In our framework, the mainstream bank is unable to update its misinformed prior in the absence of lending activity; consequently, the group remains locked into unfavorable borrowing terms.
Our work also relates to a literature on delayed feedback with respect to implementation decisions in machine learning settings Liu et al. [2018], Pagan et al. [2023], Chaney et al. [2018]. We describe a setting in which mainstream banks are stuck in a negative feedback loop: their failure to learn about applicant creditworthiness is self sustaining.
@@
Finally, our work contributes to a broader effort to model rational but suboptimal decision-making by or with respect to disadvantaged groups.Diana et al. [2024] models pessimism traps and herding behavior as individuals grapple with cycles of noisy and censored information when making decisions about potentially higher reward but riskier ends. Our argument is similar in broad outline to that of FOSTER and VOHRA [1992], Hu and Chen [2018], who argue that temporary interventions in the labor market to address discrimination may lead to long-run improvements in the fairness of observed decisions.
2 Preliminaries and Model
@@
We analyze a multi-period environment indexed by $t=1,2,\ldots,T$, where $T$ may be arbitrarily large or unbounded. At the onset of each period $t$, the high-interest (subprime) bank selects an interest-rate premium $\nu^{(t)}$. Each loan applicant belongs to one of two groups (see Section 2.1) and applies simultaneously to both the mainstream (low-interest) bank and the subprime bank. The mainstream bank charges a baseline rate normalized to $1$, while the subprime bank charges $1+\nu^{(t)}$.
@@
Each bank evaluates whether to approve or reject each applicant in that period, seeking to maximize its expected profit subject to a risk-management constraint (detailed in Section 2.2.2). If an applicant receives approvals from both banks, the applicant typically accepts the offer with the lower interest rate; if only one bank approves the loan, that solitary offer is the only available option. After the loans are finalized, the realized payoff from each executed loan is observed exclusively by the bank that issued it. Consequently, a bank that rejects an application or whose offer is declined does not observe the repayment outcome for that applicant.
@@
At the end of period $t$, both banks update their beliefs regarding the repayment distribution for each group, employing a Bayesian learning procedure described in Section 2.2.4. These updated beliefs may affect the subprime bank’s subsequent choice of premium $\nu^{(t+1)}$ and both banks’ lending decisions in period $t+1$. This iterative process continues until period $T$, yielding a dynamic model in which information acquisition and evolving risk assessments influence lending behavior over time.
2.1 Loan Applicants
@@
There are two groups, denoted by $i\in\{W,B\}$. $W$ represents the racial majority, while $B$ represents the racial minority. Each group has a payoff distribution $\pi_{i}^{(t)}\sim\mathcal{N}(\mu_{i},\sigma_{i}^{2})$. The bank’s return from appproving a loan depends on this realized payoff.
@@
$\forall t:\mathbb{E}\left[\pi^{(t)}_{W}\right]=\mu_{W}=\mu_{B}=\mathbb{E}\left[\pi^{(t)}_{B}\right]>0$
@@
First, we assume that both groups have the same expected creditworthiness. We do this because we want to study the specific case where minority and majority racial groups are equally creditworthy. Our goal is to show that discrimination is still possible in such a setting, which coincides with the stylized facts presented in Section 1. We allow their variances, $\sigma_{W}^{2}$ and $\sigma_{B}^{2}$, to differ.
@@
We suppose that each group strictly prefers to accept a loan rather than not, and strictly prefers to accept the cheapest loan. We assume that applicants randomize if both loans have the same price. We denote each group’s decision to accept the loan with $S_{i}\in\{0,1\}$
2.2 Lenders
@@
The model features two banks: Bank $L$, which offers low-interest (mainstream) loans, and Bank $H$, which provides high-interest (subprime) loans. We refer to banks with the index $j$. In each period, both banks decide whether to extend a loan to each group. Let $\bm{A}_{j}^{(t)}\in\{0,1\}^{2}$ denote the vector of approval decisions made by the bank $j$ at time $t$.
2.2.1 Bank payoff functions
@@
If the bank issues a loan to an applicant from group $i$ in period $t$, it earns a payoff that depends on the realized value of $\pi_{i}^{(t)}$. We normalize the interest rate of Bank $L$ to $1$, so that its profit $\Pi$ in period $t$ is given by
$\Pi_{L}^{(t)}=\sum_{i}S_{i}^{(t)}A_{iL}^{(t)}\pi_{i}^{(t)}$
In contrast, Bank $H$ charges an interest rate of $1+\nu$, with $\nu>0$. It has profit function:
$\Pi_{H}^{(t)}=\sum_{i}S_{i}^{(t)}A_{iH}^{(t)}(1+\nu^{(t)})\left[\max\left\{\pi_{i}^{(t)},0\right\}+\min\left\{\pi_{i}^{(t)},0\right\}\right].$
@@
Here, positive returns are amplified by a factor of $1+\nu^{(t)}$, whereas losses are incurred on a dollar-for-dollar basis.
2.2.2 Risk Management Constraints
@@
We next suppose that banks face a risk management or solvency constraint. That is, there is some level of financial loss that is ‘unacceptable’ to the bank. This may be due to regulatory constraints, such as Basel III or Dodd-Frank, or liquidity constraints Li [2014]. This loss threshold is probabilistic: the bank is willing to accept some nonzero risk that their loss from making a loan falls below a certain threshold, but wants to specifically limit the probability that this occurs to less than a risk tolerance $\alpha\%$. Typically, this value is set to $1\%$ or $5\%$Jorion [2000].
@@
We model this with a per-period Value-at-Risk ($\mathrm{VaR}$) constraint Artzner et al. [1999]. That is, we require that the bank’s anticipated profit in each period must be greater than some constant $\rho_{j}<0$ with probability at least $1-\alpha$. We can think of $\rho_{j}$ as the bank’s maximum acceptable loss. Equivalently, the bank is willing to accept an $\alpha\%$ risk that their profit will fall below the bank’s risk management threshold $\rho_{j}$.
@@
$\mathbb{P}\!\left(\Pi_{j}^{(t)}<\rho_{j}\right)\leq\alpha$
Which is simply that the bank will accept a risk of at most $\alpha\%$ that their profit in period $t$ falls below $\rho_{j}$.
@@
$\underset{\bm{A}_{j}^{(t)}}{\arg\max}\;\;\Pi_{j}^{(t)}\left(\bm{A}_{j}^{(t)}\right)\quad\text{subject to}\quad\mathbb{P}\!\left(\Pi_{j}^{(t)}\left(\bm{A}_{j}^{(t)}\right)\leq\rho_{j}\right)\leq\alpha$
@@
2.2.3 Bank beliefs
The actual profit $\Pi_{j}^{(t)}$ is unknown ex ante, however: it is only observed once a loan has been approved and accepted. Banks must therefore instead make decisions based on prior beliefs about the profitability of approving a loan. Denote by $\hat{\Pi}_{j}^{(t)}$ the bank’s posterior estimate of profit from lending at time $t$. Each bank instead must solve the feasible problem:
$\underset{\bm{A}_{j}^{(t)}}{\arg\max}\;\;\hat{\Pi}_{j}^{(t)}\left(\bm{A}_{j}^{(t)}\right)\quad\text{subject to}\quad\mathbb{P}\!\left(\hat{\Pi}_{j}^{(t)}\left(\bm{A}_{j}^{(t)}\right)\leq\rho_{j}\right)\leq\alpha$
Which, importantly, depends on bank beliefs about applicants. We assume that each bank $j$ has initial prior beliefs regarding the parameters of the distribution of returns from issuing a loan to a member of group $i$, specifically:
@@ 2.2.3 Bank beliefs
Each bank $j$ has prior beliefs:
$\left\{\hat{\mu}_{ij}^{(0)},\hat{\sigma}_{ij}^{(0)}\right\}_{i\in\{W,B\}}$
We assume that both banks have accurate beliefs about the variance of group $W$. This does not have significant implications for the analysis, since the problem of interest is how banks respond to differentially imperfect information about lenders’ creditworthiness. This assumption also reflects empirical findings that credit score data is differentially missing for minority group applicants versus majority group applicants Kenneth P. Brevoort [2015].
Assumption 2 (Group W’s variance is common knowledge).
We assume that ${\hat{\sigma}_{W}}^{(0)}=\sigma_{W}$ for both banks.
2.2.4 Learning through lending

### v2

@@ Abstract
Disparities in lending to minority applicants persist even as algorithmic lending finds widespread adoption. We study the role of risk-management constraints, specifically Value-at-Risk ($\mathrm{VaR}$) and Expected Shortfall (ES), in inducing inequality in loan approval decisions, even among applicants who are equally creditworthy. Empirical research finds that disparities in the interest rates charged to minority groups can remain large even when loan applicants from different groups are equally creditworthy. We contribute an original analysis of 431,551 loan applications recorded under the Home Mortgage Disclosure Act, illustrating that disparities in data quality are associated with higher rates of loan denial and higher interest rate spreads for Black borrowers. We develop a formal model in which a mainstream bank (low-interest) is more sensitive to variance risk than a subprime bank (high-interest). If the mainstream bank has an inflated prior belief about the variance of the minority group, it may deny that group credit indefinitely, thus never learning the true risk of lending to that group, while the subprime lender serves this population at higher rates. We call this “The Subprime Trap”: an equilibrium in which minority lenders can borrow only from high-cost lenders, even when they are as creditworthy as majority applicants. Finally, we show that a finite subsidy can help minority groups escape the trap: subsidies cover enough of the mainstream bank’s downside risk so that it can afford to lend to, and thereby learn the true risk of lending to, the minority group. Once the mainstream bank has observed sufficiently many loans, its beliefs converge to the true underlying risk, it approves the applications of minority groups, and competition drives down the interest rates of subprime loans.
Introduction
@@
Algorithmic lending has grown rapidly as scalable ML methods achieve wide adoption among both existing lenders and new market entrants, bringing with them the possibility of significantly improving the fairness of financial decisions based on observable data about loan applicants (Khandani, Kim, and Lo 2010; Bono, Croxson, and Giles 2021; Berg, Fuster, and Puri 2022; Remolina 2022). However, inequalities persist in both loan approval rates and interest rates charged to minority applicants versus white applicants, and a switch to algorithmic lending procedures does not necessarily improve outcomes on either metric (Romei and Ruggieri 2013; Žliobaitė 2017; Quillian, Lee, and Honoré 2020; Giacoletti, Heimer, and Yu 2021; Aliprantis, Carroll, and Young 2022; Bartlett et al. 2022; Fuster et al. 2021). These inequities exist against a background of historic discrimination in US retail banking to individuals (Quillian, Lee, and Honoré 2020), businesses (Blanchflower, Levine, and Zimmerman 2003), and a long-standing “racial wealth gap”: persistent differences in median household wealth by ethnic group (Charles and Hurst 2002; Derenoncourt et al. 2023; Althoff and Reichardt 2024).
@@
Notably, disparities in interest rates and loan approvals persist even when minority applicants have comparable credit scores to majority applicants (Bayer, Ferreira, and Ross 2016; Popick 2022; Crosignani and Le 2023). This combination of facts creates a puzzle: are lenders selecting on observable risk factors? And if not, why, besides explicit discrimination, might they be failing to lend to minority applicants?
@@
Several empirical findings help us to understand this puzzle. First, ethnic groups sort across lenders, with minorities more likely to accept loans from high-interest rate banks versus conventional lenders, even conditional on credit score. (Bayer, Ferreira, and Ross 2016) write:
@@
Modern algorithmic lending systems operationalize risk assessment through machine learning models trained on historical data. Minority applicants disproportionately have “thin credit files” with limited payment history data, and this data sparsity directly manifests as higher prediction variance in ML models (Blattner and Nelson 2021a). When banks face risk-management constraints, the higher predicted variance translates into higher rates of loan denial. Because denied applicants never generate repayment data, banks cannotthen learn that their variance estimates were inflated, creating a self-reinforcing equilibrium of exclusion.
@@
It is this set of stylized facts that motivates our model. How do minorities with good credit scores end up with no choice but to accept high-cost loans? To answer this question, we consider the role of risk management constraints when banks have imperfect information about their applicant pool. Lending is a problem of imperfect information: banks do not know the exact probability that a given borrower will repay a loan, and so must use observable characteristics as a proxy for repayment probability (Adams, Einav, and Levin 2009; Crawford, Pavanini, and Schivardi 2018). Banks rely on credit scoring, which is subject to known biases, and has lower quality information about the creditworthiness of minority applicants (Blattner and Nelson 2021a). Reliance on credit scores can therefore lead banks to hold inaccurate prior beliefs about minority applicants, which can sustain equilibria in which minority groups systematically lose out (Coate and Loury 1993; Kim and Loury 2018).
@@
We study the effect of Value-at-Risk ($\mathrm{VaR}$) constraints on loan approval decisions (Artzner et al. 1999; Jorion 2000). Our analysis focuses on a formal model in which a mainstream, or “low interest rate”, bank exhibits heightened sensitivity to variance risk relative to a subprime, or “high interest rate”, bank. We show that when the mainstream bank has inflated prior beliefs about the variance repayments from the minority group, $\mathrm{VaR}$ constraints bind, which means that it will then systematically refrain from extending credit to that group. This reluctance then prevents the bank from updating its beliefs regarding the true risk, thereby locking minority borrowers into higher-cost subprime lending arrangements. We describe this mechanism as “the subprime trap”.
Our Contributions
@@
First, we contribute an analysis of risk management constraints to the study of algorithmic fairness. We help address a puzzle of general interest: why might banks fail to extend loans even when they would otherwise be profitable? Second, we bring details of real-world banking practices into the study of algorithmic fairness. We introduce a two-bank equilibrium framework in which the mainstream bank enforces a more stringent $\mathrm{VaR}$ requirement than its subprime counterpart. Under these conditions, an erroneous initial estimate of the variance of the minority group leads the mainstream bank to permanently refrain from lending to that group. Third, we formalize the resulting equilibrium (1) and demonstrate that the minority group is consequently confined to subprime loans, despite having the same average creditworthiness as the majority group. Fourth, we establish that the provision of a modest, finite subsidy, or partial guarantee, can resolve this equilibrium inefficiency. By adequately mitigating downside risk, the subsidy enables the main bank to lend and learn about the creditworthiness of the minority group. With an updated assessment of risk, the bank eventually meets its $\mathrm{VaR}$ requirements without continued external intervention, thereby extending mainstream credit at favorable rates.
@@
This investigation contributes to the literature by highlighting how inaccurate risk metrics, in conjunction with risk-based capital constraints, can systematically exclude certain borrower groups from low-interest credit. Our results further imply that targeted subsidies can effectively rectify informational failures, yielding improved outcomes for both lenders and minority borrowers.
Related Literature
@@
Our work contributes to several strands of literature, including algorithmic fairness in lending, the economics of discrimination, and the role of informational frictions in sustaining suboptimal market equilibria. In the domain of algorithmic fairness, recent research has examined how increasingly sophisticated data-driven methods in credit scoring may inadvertently perpetuate or exacerbate disparities (Barocas and Selbst 2016; Barocas, Hardt, and Narayanan 2023; Kumar, Hines, and Dickerson 2022). While algorithmic approaches promise enhanced accuracy, empirical evidence indicates that minority applicants often face higher interest rates or are denied credit outright (Bartlett et al. 2022; Crosignani and Le 2023).
@@
Discrimination is a classical topic in economics, studied in a large theoretical and empirical literature (Becker 1971; Phelps 1972; Lundberg and Startz 1983; Arrow 1998; Heckman 1998; Cowgill and Tucker 2019). Building on this tradition, subsequent studies have explored how negative stereotypes or misperceptions of risk may translate into adverse outcomes for minority borrowers (Fang and Moro 2011; Coate and Loury 1993). In a manner analogous to these analyses, our work demonstrates that inflated beliefs about repayment variance can result in a self-reinforcing equilibrium whereby minority groups are persistently relegated to subprime lending markets.
A related literature examines the role of informational frictions and the corrective impact of policy interventions (Akerlof 1970; Akerlof and Shiller 2015; Thaler 2016). (Cai et al. 2020) studied contexts in which banks can acquire improved information about borrowers’ creditworthiness through selective experimentation. (Donahue and Barocas 2021) study the trade-offs between solidaristic insurance policies, in which minorities are subsidized in accordance with risks that are causally related to their minority status, and actuarially-fair insurance policies, in which everyone pays premia equal to their marginal risk.
Our work also relates to work on delayed feedback with respect to implementation decisions in machine learning settings (Liu et al. 2018; Pagan et al. 2023; Chaney, Stewart, and Engelhardt 2018). We describe a setting in which mainstream banks are stuck in a negative feedback loop: their failure to learn about applicant creditworthiness is self-sustaining. Our work is also related to literature on suboptimal outcomes with bandits. (Honda and Takemura 2013) prove that Thompson Sampling’s optimality depends critically on prior specification, while (Ghosh, Chowdhury, and Gopalan 2017) show that any algorithm achieving optimal performance under perfect model specification must suffer linear regret under misspecification. This mirrors our setting where banks with incorrect variance priors never explore lending to certain groups. The dueling bandits framework of (Yue et al. 2012) captures competition between lenders, establishing information-theoretically optimal algorithms for pairwise comparisons. (Frazier et al. 2014) study how to incentivize exploration when arms are pulled by self-interested agents – our approach, in which subsidies encourage banks to ‘learn through lending’, is closely related to this setting.
@@
Finally, our work contributes to a broader effort to model rational but suboptimal decision-making by or with respect to disadvantaged groups.(Diana et al. 2024) model pessimism traps and herding behavior as individuals grapple with cycles of noisy and censored information when making decisions about potentially higher reward but riskier ends. Our argument is similar in broad outline to that of (Foster and Vohra 1992; Hu and Chen 2018), who argue that temporary interventions in the labor market to address discrimination may lead to long-run improvements in the fairness of observed decisions.
Empirical Motivation
We analyzed 2024 Home Mortgage Disclosure Act (HMDA) data, examining 431,551 conventional home purchase mortgage applications by single applicants listed as Black or White on loan applications. The HMDA requires that financial institutions that report at least 60,000 applications and loans in the previous calendar year must report data on lending applicants and loan outcomes. We examine racial disparities in two lending outcomes: loan denial decisions ($N=431,551$) and interest rate spreads conditional on approval ($N=247,329$).
Table 1: Lending Outcomes by Race and Data Quality
Panel A: Denial Rates
Complete Data Missing Income Difference
Black 26.1% 57.6% +31.5pp
White 9.8% 23.1% +13.3pp
Differential +18.2pp
Panel B: Interest Rate Spreads [^1] Rate spread measures how much more expensive a borrower’s loan is compared to the best rates available to prime borrowers at the time of origination. A spread of 1% means the borrower pays 1 percentage point more in annual interest than the benchmark rate. On Approved Loans
Complete Data Missing Income Difference
Black 0.85% 1.12% +0.27pp
White 0.62% 0.74% +0.12pp
Differential +0.15pp
This analysis highlights that Black applicants with missing income data face higher loan denial rates both marginally, and conditional on having missing data; and, if approved, pay higher interest rate spreads both marginally, and conditional on having missing income data.
While this analysis is descriptive, it suggests a puzzle: why do differences in data quality lead to higher rejection rates and higher interest rates for Black applicants specifically? We propose an explanation for this observed phenomenon in our formal model below.
Preliminaries and Model
@@
We analyze a multi-period environment indexed by $t=1,2,\ldots,T$. At the onset of each period $t$, the high-interest (subprime) bank selects an interest-rate premium $\nu_{t}\in[0,\nu^{\text{max}}]$. Each loan applicant belongs to one of two groups and applies simultaneously to both the mainstream (low-interest) bank and the subprime bank. The mainstream bank charges a baseline rate normalized to $1$, while the subprime bank charges $1+\nu_{t}$.
@@
Each bank evaluates whether to approve or reject each applicant in that period, seeking to maximize its expected profit subject to a risk-management constraint. If an applicant receives approvals from both banks, the applicant typically accepts the offer with the lower interest rate; if only one bank approves the loan, it is accepted. After the loans are finalized, each bank privately observes its payoff. Consequently, a bank that rejects an application or whose offer is declined does not observe the repayment outcome for that applicant.
@@
At the end of period $t$, both banks update their beliefs regarding the repayment distribution for each group, employing a Bayesian learning procedure described in Section Learning Through Lending. These updated beliefs affect the subprime bank’s subsequent choice of premium $\nu_{t+1}$ and both banks’ lending decisions in period $t+1$.
Loan Applicants
Payoff distributions
@@
There are two groups, denoted by $i\in\{W,B\}$. $W$ represents the racial majority, while $B$ represents the racial minority. Each group has a payoff distribution $\pi_{it}\sim\mathcal{N}(\mu_{i},\sigma^{2}_{i})$. The bank’s return from approving a loan depends on this realized payoff.
@@
$\forall t:\mathbb{E}\left[\pi_{Wt}\right]=\mu_{W}=\mu_{B}=\mathbb{E}\left[\pi_{Bt}\right]>0$
@@
First, we assume that both groups have the same expected creditworthiness. We do this because we want to study the specific case where minority and majority racial groups are equally creditworthy. We allow their variances, $\sigma_{W}^{2}$ and $\sigma_{B}^{2}$, to differ.
Second, we assume that this is common knowledge: both banks know that each group is equally creditworthy.
@@
We suppose that each group strictly prefers to accept a loan rather than not, and strictly prefers to accept the cheapest loan. We assume that applicants randomize if both loans have the same price. We denote each group’s decision to accept the loan with $S_{i}\in\{0,1\}$.
Observable Applicant Characteristics
We suppose that individuals are associated with repayment-probability-relevant applicant characteristics: repayment history, assets, income, net debt, and so forth. We suppose that this information is differentially missing by group: minority lenders have fewer complete characteristics. We formalize this simply by supposing that for each group, the number of complete applicant files, or the credit history for group $i$, $C_{i}$ is binomial, so that we observe $np_{i}$ complete applicant files and $n(1-p_{i})$ incomplete applicant files.
Assumption 2 (Differential Information about Applicants).
$p_{B}<p_{W}\text{ so that }\mathbb{E}[C_{B}]<\mathbb{E}[C_{W}]$
For simplicity, we set $p_{W}=1$, and $p_{B}<1$. This is WLOG, because what matters for the derivation of our results is the difference in information between the groups.
Lenders
@@
The model features two banks: Bank $L$, which offers low-interest (mainstream) loans, and Bank $H$, which provides high-interest (subprime) loans. We refer to banks with the index $j$. In each period, both banks decide whether to extend a loan to each group. $A_{ijt}\in\{0,1\}$ denotes whether bank $j$ approves a loan to group $i$ at time $t$.
Bank payoff functions
@@
If the bank issues a loan to an applicant from group $i$ in period $t$, it earns the payoff $\pi_{it}$. We normalize the interest rate of Bank $L$ to $1$, so that its profit $\Pi$ in period $t$ is given by
$\Pi_{Lt}=\sum_{i}S_{iLt}A_{iLt}\pi_{it}$
Bank $H$ charges an interest rate of $1+\nu_{t}$. It has profit function:
$\Pi_{Ht}=\sum_{i}S_{iHt}A_{iHt}^{(t)}\left[(1+\nu_{t})\max\left\{\pi_{it},0\right\}+\min\left\{\pi_{it},0\right\}\right].$
@@
Here, positive returns are amplified by a factor of $1+\nu_{t}$, whereas losses are incurred on a dollar-for-dollar basis.
Risk Management Constraints
@@
We next suppose that banks face a risk management or solvency constraint. That is, there is some level of financial loss that is ‘unacceptable’ to the bank. This may be due to regulatory constraints, such as Basel III or Dodd-Frank, or liquidity constraints (Li 2014). This loss threshold is probabilistic: the bank is willing to accept some nonzero risk that their loss from making a loan falls below a certain threshold, but wants to specifically limit the probability that this occurs to less than a risk tolerance $\alpha\%$. Typically, this value is set to $1\%$ or $5\%$ (Jorion 2000).
@@
We model this with a per-period Value-at-Risk ($\mathrm{VaR}$) constraint (Artzner et al. 1999). That is, we require that the bank’s anticipated profit in each period must be greater than some constant $\rho<0$ with probability at least $1-\alpha$. We can think of $\rho$ as the bank’s maximum acceptable loss. Equivalently, the bank is willing to accept an $\alpha\%$ risk that their profit will fall below the bank’s risk management threshold $\rho$.
@@
$\mathbb{P}\!\left(\Pi_{jt}<\rho\right)\leq\alpha$
This simply states that the bank will accept a risk of at most $\alpha\%$ that their profit in period $t$ falls below $\rho$.
@@
$\underset{\bm{A}_{jt}}{\arg\max}\;\;\Pi_{jt}\left(\bm{A}_{jt}\right)\quad\text{subject to}\quad\mathbb{P}\!\left(\Pi_{jt}\left(\bm{A}_{jt}\right)\leq\rho\right)\leq\alpha$
@@
Prior Bank Beliefs
The actual profit $\Pi_{jt}$ is unknown ex ante, however: it is only observed once a loan has been approved and accepted. Banks must therefore instead make decisions based on prior beliefs about the profitability of approving a loan.
Note that, at $t=1$, the bank cannot use observed repayment information in the determination of the loan application. Instead, we suppose that the bank uses historical credit data to estimate the risk of lending to specific groups.
Denote by $\hat{\Pi}_{jt}$ the bank’s posterior estimate of profit from lending at time $t$. Each bank instead must solve the feasible problem:
$\underset{\bm{A}_{jt}}{\arg\max}\;\;\hat{\Pi}_{jt}\left(\bm{A}_{jt}\right)\quad\text{s. t.}\quad\mathbb{P}\!\left(\hat{\Pi}_{jt}\left(\bm{A}_{jt}\right)\leq\rho\right)\leq\alpha$
Which, importantly, depends on bank beliefs about applicants.
We assume that each bank $j$ has initial prior beliefs regarding the parameters of the distribution of returns from issuing a loan to a member of group $i$.
We assume, as before, that the mean of both payoff distributions is known and common knowledge. We assume that the scale of each distribution is unknown, however.
@@
We suppose that each bank has an Inverse-Gamma prior on the variance of each group, so that:
$\sigma^{2}_{ij,0}\sim\text{Inv-}\Gamma(a_{0},b_{0})$
Differential Credit Records Drive Prior Variances
We suppose that each bank estimates a model of repayment probability based on historical credit data to form its initial beliefs. That is, on having observed a credit file with effective sample size $C_{i}=np_{i}$, each bank forms a prior about each group’s variance as follows:
$\sigma^{2}_{ij}|C_{i}\sim\text{Inv-}\Gamma\left(a_{0}+\dfrac{np_{i}}{2},b_{0}+\dfrac{np_{i}S^{2}_{i}}{2}\right)$
Since we suppose that $p_{B}<p_{W}=1$, we have (assuming that $S^{2}_{B}=S^{2}_{W}=1$):
$\sigma^{2}_{Bj,0}|C_{B}$ $\sim\text{Inv-}\Gamma\left(a_{0}+\dfrac{np_{B}}{2},b_{0}+\dfrac{np_{B}}{2}\right)$
$\sigma^{2}_{Wj,0}|C_{W}$ $\sim\text{Inv-}\Gamma\left(a_{0}+\dfrac{n}{2},b_{0}+\dfrac{n}{2}\right)$
So that:
$\hat{\sigma}^{2}_{Bj,0}=\mathbb{E}[\sigma^{2}_{Bj,0}|C_{B}]>\mathbb{E}[\sigma^{2}_{Wj,0}|C_{W}]=\hat{\sigma}^{2}_{Wj,0}$
Or, in other words, based on observable credit histories, each bank has a prior belief that the variance of the minority group is larger than the variance of the majority group.
Learning Through Lending